If ( x*2 +1/x)*2 =15



Then x=1/x=?


1
Expert's answer
2021-12-30T16:13:29-0500

"(2x+\frac{1}{x})*2=15".

"4x+\frac{2}{x}=15."

"4x^2-15x+2=0."

"x_1=\frac{15-\sqrt{193}}{8}\approx 0.1384."

"x_2=\frac{15+\sqrt{193}}{8}\approx 3.6116."